OpenStudy (anonymous):

Solve: 2(x + 6) − (x + 1) = 16 3 5 2 1

OpenStudy (anonymous):

5

OpenStudy (anonymous):

its simple.........

OpenStudy (anonymous):

cos 2x+12-x-1=16 so x=16-11 so x=5

OpenStudy (anonymous):

use the distributive property 2(x + 6) − (x + 1) = 16 2x + 12 -x -1 = 16 combine like terms 2x -x + 12 - 1 = 16 x + 11 = 16 then subtract both sides by 11 x = 5